the leght of a rectangle is 5 cm greater than its width. the perimeter is 58 cm. find the dimension of the rectangle
Found 2 solutions by checkley77, rapaljer:
Answer by checkley77(12844)

L=W+5
2L+2W=58
2(W+5)+2W=58
2W+10+2W=58
4W=58-10
4W=48
W=48/4
W=12 cm.
L=12+5=17 cm.
PROOF:
2*17+2*12=58
34+24=58
58=58

Answer by rapaljer(4671)

Let x = width of the rectangle
x+5 = length

2W + 2L = P

2(x) + 2(x+5) = 58
2x+2x+10 = 58
4x+10=58
4x = 48
x=12 width
x+5 = 12+5=17 length

Check:
2*12+2*17
24+34= 58 = Perimeter

For additional explanation on solving Word Problems please see my website by clicking on my tutor name "Rapaljer" anywhere in algebra.com. On my homepage, look for the link "Basic, Intermediate, and College Algebra: One Step at a Time." Go to "Basic Algebra", and look for "Chapter 1" and select the sections marked "Word Problems." See also my MATH IN LIVING COLOR pages, where a LOT of these problems are solved in COLOR--it really makes a difference!! My students have ALL told me that my explanations are easier than traditional books. I WROTE ALL OF THESE SECTIONS SPECIFICALLY FOR PEOPLE WHO DON'T GET MATH!! I hope it helps you!!

R^2

Dr. Robert J. Rapalje, Retired
Altamonte Springs Campus
Seminole Community College
Florida